# A Two-Stage Pretraining-Finetuning Framework for Treatment Effect Estimation with Unmeasured Confounding

### v1

@@ Abstract.
Estimating the conditional average treatment effect (CATE) from observational data plays a crucial role in areas such as e-commerce, healthcare, and economics. Existing studies mainly rely on the strong ignorability assumption that there are no unmeasured confounders, whose presence cannot be tested from observational data and can invalidate any causal conclusion. In contrast, data collected from randomized controlled trials (RCT) do not suffer from confounding, but are usually limited by a small sample size. In this paper, we propose a two-stage pretraining-finetuning (TSPF) framework using both large-scale observational data and small-scale RCT data to estimate the CATE in the presence of unmeasured confounding. In the first stage, a foundational representation of covariates is trained to estimate counterfactual outcomes through large-scale observational data. In the second stage, we propose to train an augmented representation of the covariates, which is concatenated to the foundational representation obtained in the first stage to adjust for the unmeasured confounding. To avoid overfitting caused by the small-scale RCT data in the second stage, we further propose a partial parameter initialization approach, rather than training a separate network. The superiority of our approach is validated on two public datasets with extensive experiments. The code is available at https://github.com/zhouchuanCN/KDD25-TSPF.
Causal Effect Estimation, Unmeasured Confounding, Data Fusion
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709161isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
The conditional average treatment effect (CATE) is the average causal effect of a treatment or an intervention on the outcome of interest given the covariates (Pearl, 2009), which plays an important role in diverse fields, such as e-commerce (Wu et al., 2022b), healthcare (Robins and Hernán, 2016), and economics (Huynh et al., 2016). In e-commerce, the platforms desire to predict how recommending a specific product to a particular user affects the probability of purchase (Li et al., 2022), and thereby influence the total profit. In healthcare, doctors assess the potential outcome for different patient groups when administering a certain treatment (Corrao et al., 2012) for precision medicine. Similarly in economics, the policymakers evaluate how much a job training program will raise employment opportunities for unemployed individuals (Britto et al., 2022).
@@
To enhance the accuracy of CATE estimation, representation-based learning approaches have gathered increasing attention due to their impressive performance (Johansson et al., 2016; Li et al., 2024a; Zhu et al., 2024a; Wang et al., 2024). These approaches focus on generating covariate representations, with the objective of mitigating confounding bias by minimizing distributional discrepancies between the treatment group and control group. To obtain such representations, previous approaches have developed substantial theory and explored extensive practice. For instance, some of them use integral probability metric (IPM) for regularization (Johansson et al., 2016), while a few approaches emphasize local similarity preservation (Yao et al., 2018), targeted learning (Zhang et al., 2020), and optimal transport (Wang et al., 2024).
@@
However, the aforementioned methods may ignore unmeasured confounding, which is very common in real-world scenarios. In our e-commerce example, the financial status of users should be sensitive and cannot be collected (Li et al., 2023). In the healthcare case, the personal lifestyles of patients are difficult to obtain (Charpignon et al., 2022). For the example in economics, personal working status is difficult to measure (Xu and Li, 2020). These unmeasured variables can affect the treatment and outcome simultaneously, which causes confounding bias. Therefore, proposing methods to account for the confounding bias is crucial to accurately estimate CATE.
@@
To address unmeasured confounding, one category of mainstream methods can only rely on large-scale observational (OBS) data, including sensitivity analysis (Imbens, 2003; Dorn et al., 2024), front door adjustment methods (Fulcher et al., 2020; Zhu et al., 2024b), and instrumental variables methods (Angrist et al., 1996; Rudolph et al., 2024). These methods require additional strong assumptions that cannot be tested from the data and raise concerns if these assumptions are violated (Hartwig et al., 2023; Kong et al., 2025). Compared to the OBS data, randomized controlled trial (RCT) data are considered as the gold standard for causal effect estimation (Prosperi et al., 2020). However, practical challenges such as high costs and ethical concerns may make the collection of RCT data difficult (Zabor et al., 2020; Bédécarrats et al., 2020), resulting in limited sample sizes. Due to the small size, it is impractical to directly train causal effect prediction models on RCT data alone (Hoogland et al., 2021). Therefore, it is necessary to find an effective method to combine small-scale RCT data with large-scale OBS data. Some studies use methods of correcting for residuals, in which RCT data are utilized to correct for biased estimates from OBS data (Colnet et al., 2024). However, these methods still rely on strong assumptions, including linear and additive data generation processes (Kallus et al., 2018) and a shared structure between the two data types (RCT and OBS) (Hatt et al., 2022b). This can lead to poor model performance in the complex real world. In contrast, other studies without these assumptions generally use both types of data to estimate CATE and the residuals between the biased and unbiased estimates (Wu and Yang, 2022), yet these methods suffer from the risk of overfitting due to insufficient RCT sample sizes.
@@
In this paper, we introduce a two-stage framework named TSPF for CATE estimation with unmeasured confounding based on the pretraining-finetuning principle. Our approach leverages large-scale OBS data to train a foundational representation of covariates and then uses relatively small-scale RCT data to adjust the representation learned from OBS data. We then train a more accurate prediction model using this adjusted representation. In the second stage, we introduce an additional module that ensures stronger representation ability compared to the methods that use RCT data to estimate the residuals between the biased and unbiased estimates.
@@
  • We present a two-stage pretraining-finetuning framework for estimating the CATE. This framework tackles the issue of unmeasured confounding by using a small amount of unconfounded RCT data to calibrate the representations learned from observational (OBS) data.
@@
  • Extensive experiments conducted on the IHDP and Jobs datasets demonstrate the effectiveness of our approach.
2. Related Work
2.1. CATE Estimation
CATE also known as heterogeneous treatment effect (HTE), refers to the average causal effects of a treatment on an outcome for subgroups with different covariates. To accurately estimate CATE, early statistical methods include matching (Dehejia and Wahba, 2002), stratification (O’Muircheartaigh and Hedges, 2014), reweighting (Rosenbaum, 1987; Bang and Robins, 2005), and tree-based methods (Chipman et al., 2010; Wager and Athey, 2018). With the impressive advances in deep learning, recent CATE estimation methods can be broadly categorized into representational learning-based and generative model-based methods. Specifically, the core idea of representation learning-based methods is to learn a balanced covariate representation that has similar distributions in the treatment and control groups to mitigate confounding bias (Assaad et al., 2021; Yao et al., 2021; Guo et al., 2020). Such balanced representations can be obtained by using integral probability metric (IPM) for regularization (Johansson et al., 2016; Shalit et al., 2017), local similarity preservation (Yao et al., 2018, 2019), targeted learning (Shi et al., 2019; Zhang et al., 2020), and optimal transport (Wang et al., 2024; Torous et al., 2021). An alternative category is generative model-based methods that estimate the counterfactual outcomes by assuming their data generation process and exploiting generative models (Zhang et al., 2021; Zou et al., 2020; Guo et al., 2020). Specifically, CEVAE uses variational autoencoder (VAE) to learn confounders from observed variables (Louizos et al., 2017). SCIGAN generates missing counterfactual results based on generative adversarial networks (GAN) and combines the facts and estimated counterfactual results to estimate CATE (Bica et al., 2020). Unlike previous representation learning-based methods, we do not require no unmeasured confounding assumption. Instead, we learn an augmented representation to address the effects of unmeasured confounding for correcting biased estimates.
2.2. Unmeasured Confounding
Unmeasured confounding refers to a situation where there are unmeasured variables in the study that influence both the treatment and the outcome, which may lead to bias in CATE estimates (Ananth and Schisterman, 2018). To address this problem, previous studies can be broadly categorized into two types. The first only uses OBS data and mainly includes sensitivity analysis and auxiliary information methods. Sensitivity analysis aims to quantify the potential impact of unmeasured confounding on the treatment effect and to obtain a bound on the treatment effect (Rosenbaum and Rubin, 1983; Robins et al., 2000). However, these approaches assume a determined confounding mechanism of the unmeasured variables (Imbens, 2003). This assumption is too strong and can easily be violated in practice (Franks et al., 2019; Veitch and Zaveri, 2020). Auxiliary information methods in causal inference mainly include instrumental variable (IV) methods and front-door adjustment methods (Li et al., 2024c). IV methods rely on external variables to address unmeasured confounding in observational studies (Imbens, 2014; Wu et al., 2022a; Rudolph et al., 2024). However, these methods assume linearity and require unconfounded instruments, posing practical limitations (Frauen and Feuerriegel, 2022). Front door adjustment methods estimate the causal effect of a treatment on an outcome by leveraging a causal pathway, also known as the front-door criterion, which blocks all possible influences of unmeasured confounders (Bellemare et al., 2020; Fulcher et al., 2020). In general, front door adjustment may rely on knowing the true causal graph, which may not always be feasible in practice (Shah et al., 2024; Li et al., 2024b). Another type of approaches combine both OBS and RCT data to mitigate unmeasured confounding. Kallus et al. (Kallus et al., 2018) propose to use RCT data to correct the bias in CATE estimates derived from OBS data under the assumption that the unmeasured confounding bias is a linear function of the covariates (Kallus et al., 2018). Hatt et al. (Hatt et al., 2022a) extend to the non-linear case by using shared model structure between OBS and RCT data. Rather than using OBS and RCT data in separate regressions, by joint modeling the unmeasured confounding bias and the CATE with OBS and RCT data simultaneously as input, Wu et al. (Wu and Yang, 2022) identify both the bias from OBS estimates and unbiased CATE estimates using the R-learner approach, and Cheng et al. (Cheng and Cai, 2021) perform a weighted average of the estimates from the two data sources to mitigate the bias. Both the determination of weights and the estimation of CATE require a large sample size. However, this requirement is at odds with the small sample size of RCT. In this paper, we adopt a more flexible network architecture, with a deep neural network for predicting biased outcomes using the OBS data and extra neural networks for adjusting the representation as well as the prediction using the RCT data, which allows us to utilize limited RCT data to achieve better performance.
@@ 3.2. Identification of CATE
(Positivity) $10<e(x,G)<$ 1 for all $x\in\mathcal{X}$.
Assumption 1 is also known as no unmeasured confounding, which holds in the RCT by default due to the randomized treatment assignment. We can identify CATE based on RCT:
@@
$\omega(x)=\mathbb{E}\left[Y\mid T=1,X=x,G=0\right]-\mathbb{E}\left[Y\mid T=0,X=x,G=0\right].$
Note that due to unmeasured confounding, $\omega(x)\neq\tau(x)$ for any $x$. The difference between these two quantities is precisely the confounding effect, which we denote the residual function as:
$\eta(x)=\tau(x)-\omega(x).$
In each treatment group, we denote the residual functions:
$\eta_{t}(x)=\mathbb{E}[Y\mid T=1,X=x,G=0]-\mathbb{E}[Y\mid T=1,X=x,G=1],$
$\eta_{c}(x)=\mathbb{E}[Y\mid T=0,X=x,G=0]-\mathbb{E}[Y\mid T=0,X=x,G=1].$
3.3. Previous Work
To address the unmeasured confounding, previous work proposes many methods to estimate the residual function $\eta(x)$ and obtain an accurate estimation of CATE $\tau(x)$ with extra small-scale RCT data.
3.3.1. Residual Correction
This type of approaches first only uses the OBS data for initial estimation and then uses RCT data for correcting the biased estimation with a residual function (Kallus et al., 2018).
Step 1: Estimate $\omega(x)$ using the OBS data, denoted as $\hat{\omega}(x)$.
Step 2: Estimate $\eta(x)$ using the RCT data by minimizing:
$\min_{\eta}\sum_{i\in\mathcal{R}}(Y_{i}^{*}-\hat{\omega}(X_{i})-\eta(X_{i}))^{2},$
where $Y^{*}_{i}=\frac{T_{i}Y_{i}}{e(X_{i},G_{i})}-\frac{(1-T_{i})Y_{i}}{1-e(X_{i},G_{i})}$ is the pseudo-outcome.
The final estimate of $\tau(x)$ is $\hat{\omega}(x)+\hat{\eta}(x)$. To ensure a consistent estimate of $\tau(x)$, the residual $\eta(x)$ is assumed to be linear and additive. Moreover, the small-scale RCT data used in the second phase may cause overfitting of the estimated residual $\hat{\eta}(x)$.
3.3.2. Joint Learning of Residual and CATE Models
This type of approaches involves training the models for estimating $\eta(x)$ and $\tau(x)$ simultaneously using both OBS and RCT data (Wu and Yang, 2022).
Step 1: Estimate the conditional average outcome $\mu(X,G)$ and propensity score $e(X,G)$ by using the combined OBS and RCT data, denoted as $\hat{\mu}(X,G)$ and $\hat{e}(X,G)$.
Step 2: Estimate $\{\eta(\cdot),\tau(\cdot)\}$ by the optimization problem:
$\arg\min_{\eta,\tau}\sum_{i\in\mathcal{O}\cup\mathcal{R}}$$\left\{Y_{i}-\hat{\mu}\left(X_{i},G_{i}\right)-[\tau(X_{i})+(1-G_{i})\eta(X_{i})]\right.$
$\cdot\left.[T_{i}-\hat{e}(X_{i},G_{i})]\right\}^{2}+\Lambda(\tau)+\Lambda(\eta),$
where $\Lambda(\tau)$ and $\Lambda(\eta)$ are regularization terms on the complexity of the $\tau(\cdot)$ and $\eta(\cdot)$ functions.
Since RCT generally involves much smaller sample sizes compared to OBS studies, the data available for explicitly learning $\eta(x)$ is limited. Insufficient RCT data can lead to overfitting and a high variance of $\hat{\eta}(x)$, thus reducing the accuracy of CATE estimation.
Figure 1. The framework of our proposed method, which is composed of the modules for the first stage (blue) and second stage (yellow). Note that the two $\phi^{U}$ shown in the figure represent the same module.
@@ 4. Methodology
In this section, we present a two-stage framework for CATE estimation based on the pretraining-finetuning principle, as shown in Figure 1. The motivation is to use large-scale OBS data to train a base representation of covariates, then use relatively small-scale unbiased RCT data to calibrate the bias in the representation learned from OBS data for training the unbiased prediction model. Specifically, in the first stage, only the OBS data is used. We start with a representation module $\phi$, followed by one reconstruction module $\psi$ and two prediction heads $h_{0}$ and $h_{1}$ to ensure the learned covariate representation can have enough information and predict the outcome for control group and treatment group simultaneously. While in the second stage, we use only the RCT data. The representation module $\phi$ learned in the first stage is frozen, with a learnable representation adapter module $\phi^{U}$ to calibrate the bias of $\phi$. Then the concatenated representation is fed to two prediction heads $g_{0}$ and $g_{1}$ to obtain the unbiased predicted potential outcomes under control and treatment groups respectively. We carefully design an initialization strategy to ensure that the initialized second-stage model produces the same predictions as the converged model in the first stage. Our approach distinguishes from the one proposed by Kallus et al. (Kallus et al., 2018), which only uses linear regression to estimate the residual function $\eta(x)$ in the second stage. We can regard the first stage as pretraining on the large OBS data and the second stage as finetuning on the small-scale unbiased RCT data.
@@ 4.1.1. Representation
We design a multi-layer feed-forward neural network $\phi$ to obtain a representation $Z$ for the covariates $X$ for both treatment and control groups. In other words, for an individual sample $(X_{i},T_{i},Y_{i},G_{i}=0)$, the representation $Z_{i}=\phi(X_{i})$ remains the same whether $T_{i}=0$ or $T_{i}=1$. To better estimate the causal effect, we adopt the idea of covariate balancing on the representation space as Shalit et al. (Shalit et al., 2017) Specifically, the representations for treatment group $\{Z_{i}=\phi(X_{i}):G_{i}=0,T_{i}=1\}$ are regarded as i.i.d samples drawn randomly from a distribution $P_{\phi}^{t=1}$ and similarly $P_{\phi}^{t=0}$ for the control group. We anticipate the distributions of representations to be similar between the treatment and control groups. An integral probability metric (IPM) is employed to measure the distance between the two distributions. Thus the covariate imbalancing loss is defined as:
$\mathcal{L}_{imb}=\mathrm{IPM}_{\mathcal{G}}(\hat{P}_{\phi}^{t=0},\hat{P}_{\phi}^{t=1}),$
where $\mathrm{IPM}_{\mathcal{G}}(\cdot)$ is the empirical IPM defined by the function family $\mathcal{G}$, and $\hat{P}_{\phi}^{t=1}$ and $\hat{P}_{\phi}^{t=0}$ are empirical distributions of $P_{\phi}^{t=1}$ and $P_{\phi}^{t=0}$ respectively. In the implementation, we adopt Wasserstein distance as a showcase, which can be consistently estimated from finite samples within a mini-batch (Frogner et al., 2015).

### v2

# A Partial Initialization Strategy to Mitigate the Overfitting Problem in CATE Estimation with Hidden Confounding

@@ Abstract.
Estimating the conditional average treatment effect (CATE) from observational data plays a crucial role in areas such as e-commerce, healthcare, and economics. Existing studies mainly rely on the strong ignorability assumption that there are no hidden confounders, whose existence cannot be tested from observational data and can invalidate any causal conclusion. In contrast, data collected from randomized controlled trials (RCT) do not suffer from confounding but are usually limited by a small sample size. To avoid overfitting caused by the small-scale RCT data, we propose a novel two-stage pretraining-finetuning (TSPF) framework with a partial parameter initialization strategy to estimate the CATE in the presence of hidden confounding. In the first stage, a foundational representation of covariates is trained to estimate counterfactual outcomes through large-scale observational data. In the second stage, we propose to train an augmented representation of the covariates, which is concatenated with the foundational representation obtained in the first stage to adjust for the hidden confounding. Rather than training a separate network from scratch, part of the prediction heads are initialized from the first stage. The superiority of our approach is validated on two datasets with extensive experiments.
Causal Effect Estimation, hidden Confounding, Data Fusion
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Causal Inference and Machine Learning in Practice; August 2024; Barcelona, Spainbooktitle: CMLKDD2024: KDD 2024 Workshop, Aug 25–29, 2024, Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
The conditional average treatment effect (CATE) is the average causal effect of a treatment or an intervention on the outcome of interest given the covariates (Pearl, 2009), which plays an important role in diverse fields, such as trustworthy artificial intelligence (Li et al., 2023b), electronic commerce (Wu et al., 2022), healthcare (Robins and Hernán, 2016), and economics (Huynh et al., 2016). In e-commerce, the platforms desire to predict how recommending a specific product to a particular user affects the probability of purchase (Li et al., 2022), and thereby influence the total profit. In healthcare, doctors assess the potential outcome for different patient groups when receiving a certain treatment (Corrao et al., 2012) for precision medicine. Similarly in economics, the policymakers evaluate how much a job training program will raise employment opportunities for certain groups of unemployed individuals (Britto et al., 2022).
@@
To enhance the accuracy of CATE estimation, representation-based learning approaches have gathered increasing attention due to their impressive performance (Johansson et al., 2016; Li et al., 2024b; Zhu et al., 2024; Wang et al., 2024). These approaches focus on generating covariate representations, to mitigate confounding bias by minimizing distributional discrepancies between the treatment and control groups. Previous approaches have developed substantial theory and explored extensive practice to obtain such representations. For instance, some of them use integral probability metric (IPM) for regularization (Johansson et al., 2016), while a few approaches emphasize local similarity preservation (Yao et al., 2018), targeted learning (Zhang et al., 2020), or view the problem from new perspectives like optimal transport (Wang et al., 2024) and out-of-distribution (Li et al., 2024a).
@@
However, the aforementioned methods may ignore hidden confounding, which is very common in real-world scenarios (Li et al., 2024c). In our e-commerce example, the financial status of users might be sensitive and difficult to collect (Li et al., 2023a). In the healthcare case, the personal lifestyles of patients are difficult to obtain (Charpignon et al., 2022). For the example in economics, personal working status is difficult to measure (Xu and Li, 2020). These hidden variables can affect treatment and outcome simultaneously, which causes confounding bias in the estimation of causal effects. Therefore, proposing methods to account for the confounding bias is crucial to accurately estimate CATE.
@@
To address the hidden confounding problem, one category of mainstream methods can rely only on large-scale observational (OBS) data, including sensitivity analysis (Imbens, 2003; Dorn et al., 2024), front-door adjustment methods (Fulcher et al., 2020), and instrumental variables methods (Joshua D. Angrist and Rubin, 1996). These methods require additional strong assumptions that cannot be tested from the data and raise concerns if these assumptions are violated (Hartwig et al., 2023; Kong et al., 2025). Compared to the OBS data, randomized controlled trial (RCT) data are considered as the gold standard for causal effect estimation (Prosperi et al., 2020). However, practical challenges such as high costs and ethical concerns may make the collection of RCT data difficult (Zabor et al., 2020; Bédécarrats et al., 2020), resulting in limited sample sizes. Due to the small size, it is impractical to directly train causal effect prediction models on RCT data alone (Hoogland et al., 2021). Therefore, it is necessary to find an effective method to combine small-scale RCT data with large-scale OBS data. However, previous methods either rely on untestable strong assumptions (Kallus et al., 2018; Hatt et al., 2022b) or suffer from the risk of overfitting due to limited RCT sample sizes (Colnet et al., 2024; Wu and Yang, 2022).
@@
In this paper, we introduce a two-stage pretraining-finetuning (TSPF) framework with a partial initialization strategy for CATE estimation under hidden confounding. Our approach leverages large-scale OBS data to train a foundational representation of covariates and then uses relatively small-scale RCT data to adjust the bias in the representation learned from OBS data. We then train an unbiased prediction model using this adjusted representation. In the second stage, we introduce an additional module that ensures stronger representation ability compared to the methods that use RCT data to estimate the residuals and initialize part of the prediction heads with parameters from the first stage.
@@
  • We present the TSPF framework that includes a partial initialization strategy for CATE estimation, tackling hidden confounding by using a small amount of unconfounded RCT data to adjust the representations learned from OBS data.
@@
  • Experiments conducted on the semi-synthetic IHDP and Jobs datasets demonstrate the effectiveness of our approach.
@@ 2.2. Identification of CATE
(Positivity) $10<e(x,G=1)<$ 1.
Assumption 1 is also known as no hidden confounding, which holds in the RCT by default due to the randomized treatment assignment. We can identify CATE based on the RCT data:
@@
$\zeta(x)=\mathbb{E}\left[Y\mid T=1,X=x,G=0\right]-\mathbb{E}\left[Y\mid T=0,X=x,G=0\right].$
Note that due to hidden confounding, $\zeta(x)\neq\tau(x)$ for any $x$. The difference between these two quantities is precisely the confounding effect, which we denote the residual function as:
$\eta(x)=\tau(x)-\zeta(x).$
@@ 3. Methodology
Figure 1. The framework of our proposed method, composed of the modules for the first stage (blue) and second stage (yellow). Note that the two $\phi^{U}$ shown in the figure represent the same module.
In this section, we present a two-stage framework for the estimation of CATE based on the pretraining-finetuning principle, as shown in Figure 1. The motivation is to use large-scale OBS data to train a foundational representation of covariates, then to use relatively small-scale unbiased RCT data to remove the bias in the representation learned from OBS data for training the unbiased prediction model. Specifically, in the first stage, only the OBS data is used. We start with a representation module $\phi$, followed by one reconstruction module $\psi$ and two prediction heads $h_{0}$ and $h_{1}$ to ensure the learned covariate representation can have enough information and predict the outcome for control group and treatment group simultaneously. While in the second stage, we use only the RCT data. The representation module $\phi$ learned in the first stage is frozen, with a learnable representation adapter module $\phi^{U}$ to remove the bias of $\phi$. Then the concatenated representation is fed to two prediction heads $g_{0}$ and $g_{1}$ to obtain the unbiased predicted potential outcomes under control and treatment groups respectively. We carefully design an initialization strategy to ensure that the initialized second-stage model produces the same predictions as the converged model in the first stage. Our approach distinguishes from the one proposed by Kallus et al. (Kallus et al., 2018), which only uses linear regression to estimate the residual function $\eta(x)$ in the second stage. We can regard the first stage as pretraining on the large OBS data and the second stage as finetuning on the small-scale unbiased RCT data.
@@ 3.1.1. Representation
We design a multi-layer feed-forward neural network $\phi$ to obtain a representation $Z$ for the covariates $X$ for both treatment and control groups. In other words, for an individual sample $(X_{i},T_{i},Y_{i},G_{i}=0)$, the representation $Z_{i}=\phi(X_{i})$ remains the same whether $T_{i}=0$ or $T_{i}=1$. To better estimate the causal effect, we adopt the idea of covariate balancing on the representation space as Shalit et al. (Shalit et al., 2017). Specifically, the representations for treatment group $\{Z_{i}=\phi(X_{i}):G_{i}=0,T_{i}=1\}$ are regarded as i.i.d samples drawn randomly from a distribution $P_{\phi}^{t=1}$ and similarly $P_{\phi}^{t=0}$ for the control group. We anticipate the distributions of representations to be similar between the treatment and control groups. An integral probability metric (IPM) is employed to measure the distance between the two distributions. Thus the covariate unbalancing loss is defined as:
$\mathcal{L}_{unb}=\mathrm{IPM}_{\mathcal{G}}(\hat{P}_{\phi}^{t=1},\hat{P}_{\phi}^{t=0}),$
where $\mathrm{IPM}_{\mathcal{G}}(\cdot)$ is the empirical IPM defined by the function family $\mathcal{G}$, and $\hat{P}_{\phi}^{t=1}$ and $\hat{P}_{\phi}^{t=0}$ are empirical distributions of $P_{\phi}^{t=1}$ and $P_{\phi}^{t=0}$ respectively. In the implementation, we adopt the Wasserstein distance as a showcase, which can be consistently estimated from finite samples within a mini-batch (Frogner et al., 2015).